# Introducing Milabench: Benchmarking Accelerators for AI

### v1

@@ 2.1 Literature Review
To assist in validating GPT-4o’s predictions, the JSON template provided to the model included fields for justifications and quotes supporting each value. This strategy was applied to both research domains and model names to enhance transparency and interpretability.
The validation set was subsequently used to evaluate GPT-4o’s performance via a Multi-Label Confusion Matrix (MLCM) [11], focusing on the accuracy of predictions for research domains and model types.
@@ 2.2 Surveys
Given these biases, we relied on the expertise of the professors to suggest more pipelines than necessary but ultimately used the literature review to make the final selection.
A total of 25 pipelines were suggested, from which 10 were chosen. We reused 6 pipelines from the previous version of Milabench and added 3 additional pipelines to address obvious gaps.
@@ 3.1 Benchmarking Methodology
Where possible, datasets are generated randomly to minimize reliance on external data sources. However, in some cases, subsets of standard datasets are used when the characteristics of the data affect processing time. For example, in LLM tasks, pure random noise would result in the LLM consistently producing maximum-length sequences. The randomly generated datasets are sized to match popular real-world datasets; for instance, FakeImageNet consists of images sized at (384, 384), based on the average image sizes in ImageNet [4], and FakeVideos consists of 30 FPS, 10-second videos at (640, 480), modeled after the Kinetics dataset [13].
@@ 5 Related Work
In recent years, HPC has more widely adopted procurement practices relying strongly on application-based benchmarks in addition to pure kernel benchmarks. This transition aims to provide a more accurate assessment of system performance, aligned with the anticipated usage patterns of new systems. For example, the Jülich Supercomputer Center (JSC) has developed a benchmark suite consisting of 16 application benchmarks and 7 synthetic benchmarks [10]. Of these, 3 are AI-focused. These benchmarks include: MMoCLIP, A multi-modal text-image task using a ViT-L-14 model, trained across 8 nodes for a total of 32 GPUs; Megatron-LM, A 175 billion-parameter language model trained on a language modeling task, trained across 96 nodes for a total of 384 GPUs; and ResNet, A ResNet50 model trained for image classification, executed on 10 nodes for a total of 40 GPUs. MMoCLIP and Megatron-LM are excellent candidates to benchmark large-scale training. ResNet on the other hand uses a model that is arguably too small to benefit from large scale training on latest GPUs. The benchmark is missing important domains of AI such as RL and Graph learning. This may be intentional however as the benchmark is tailored for large scale training whereas RL and Graph learning pipelines are more generally executed on a single GPU or a single node. While MMoCLIP and Megatron-LM are well-suited to benchmark large-scale AI training, the ResNet benchmark is based on a model that may be too small to fully exploit the capabilities of modern GPUs in large-scale training scenarios. Moreover, the benchmark suite does not cover critical AI domains such as reinforcement learning (RL) and graph learning. This omission is likely intentional however, as these domains are typically executed on a single GPU or node, whereas the suite is specifically tailored for large-scale training environments.

### v2

@@ 2.1 Literature Review
To assist in validating GPT-4o’s predictions, the JSON template provided to the model included fields for justifications and quotes supporting each value. This strategy was applied to both research domains and model names to enhance transparency and interpretability. The validation set was subsequently used to evaluate GPT-4o’s performance via a Multi-Label Confusion Matrix (MLCM) [11], focusing on the accuracy of predictions for research domains and model types.
@@ 2.2 Surveys
Given these biases, we relied on the expertise of the professors to suggest more pipelines than necessary but ultimately used the literature review to make the final selection. A total of 25 pipelines were suggested, from which 10 were chosen. We reused 6 pipelines from the previous version of Milabench and added 3 additional pipelines to address obvious gaps.
@@ 3.1 Benchmarking Methodology
Where possible, datasets are generated randomly to minimize reliance on external data sources. However, in some cases, subsets of standard datasets are used when the characteristics of the data affect processing time. For example, in LLM tasks, pure random noise would result in the LLM consistently producing maximum-length sequences. The randomly generated datasets are sized to match popular real-world datasets; for instance, FakeImageNet consists of images sized at (384, 384), based on the average image sizes in ImageNet [4], and FakeVideos consists of 30 FPS, 10-second videos of size (640, 480), modeled after the Kinetics dataset [13].
@@ 5 Related Work
In recent years, HPC has more widely adopted procurement practices relying strongly on application-based benchmarks in addition to pure kernel benchmarks. This transition aims to provide a more accurate assessment of system performance, aligned with the anticipated usage patterns of new systems. For example, the Jülich Supercomputer Center (JSC) has developed a benchmark suite consisting of 16 application benchmarks and 7 synthetic benchmarks [10]. Of these, 3 are AI-focused. These benchmarks include: MMoCLIP, A multi-modal text-image task using a ViT-L-14 model, trained across 8 nodes for a total of 32 GPUs; Megatron-LM, A 175 billion-parameter language model trained on a language modeling task, trained across 96 nodes for a total of 384 GPUs; and ResNet, A ResNet50 model trained for image classification, executed on 10 nodes for a total of 40 GPUs. While MMoCLIP and Megatron-LM are excellent candidates to benchmark large-scale AI training, the ResNet benchmark is based on a model that may be too small to fully exploit the capabilities of modern GPUs in large-scale training scenarios. The benchmark suite does not cover important AI domains such as reinforcement learning (RL) and graph learning. This omission is likely intentional however, as these domains are typically executed on a single GPU or node, whereas the suite is specifically tailored for large-scale training environments.